At December 31, 2019, Whispering Corporation had the following stock outstanding.

10% cumulative preferred stock, $100 par, 109,362 shares $10,936,200
Common stock, $5 par, 4,092,240 shares 20,461,200


During 2020, Whispering did not issue any additional common stock. The following also occurred during 2020.

Income from continuing operations before taxes $22,753,600
Discontinued operations (loss before taxes) $3,254,900
Preferred dividends declared $1,093,620
Common dividends declared $2,300,800
Effective tax rate 35 %


Compute earnings per share data as it should appear in the 2020 income statement of Whispering Corporation. (Round answers to 2 decimal places, e.g. 1.48.)

Earnings Per Share

Net Income/ (Loss) $enter a dollar amount
Discontinued Operations enter a dollar amount
Income from Continuing Operations $enter a total amount
Answer #1

Answer:-

Earnings Per Share

Net Income/ (Loss) $2.83
Discontinued Operations $0.51
Income from Continuing Operations $3.34

Explanation:-

Income from continuing operations before taxes $22753600
less Taxes at 35% $7963760
Income from continuing operations after taxes $14789840

EPS: continuing operations = (Net income - preferred dividend)/No. of common shares outstanding
= ($14789840-- $1093620)/ 4092240 = $3.34

-----------------------
Discontinued operations (loss before taxes) (3254900)
less tax benefits at 35% $1139215
Discontinued operations (loss after taxes) ($2115685)

EPS: discontinued operations = ($2115685/4092240)* = ($0.51)
*for intermediary items like discontinued operations, you don't need to deduct preferred dividends.

EPS:net income

=(14789840 - 2115685 -1093620) /4092240

=2.829 ~ 2.83